OpenStudy (anonymous):

How to prove the following inequality by induction. show that 7*n^3 <5^n for all n>4 (natural numbers)

OpenStudy (anonymous):

So I proved the base case for n=4 \[7*n^3<5^n => 448 <625\] then I try to check for n +1 but get stuck so I know that \[ \text{for } n+1 we have 7(n+1)^3<5^{n+1} \] \[ 5*7n^3\le 7(n+1)^3 <5*5n^3 \] but then I am stuck to show that this is true.

OpenStudy (zarkon):

\[7(n+1)^3=7n^3+3\cdot 7n^2+3\cdot 7n+1\] \[<7n^3+n\cdot 7n^2+n\cdot 7n+1\] \[=7n^3+7n^3+ 7n^2+1\] \[<7n^3+7n^3+ 7n^3+7n^3\] \[<5^n+5^n+ 5^n+5^n=4\cdot 5^n>5\cdot5^n=5^{n+1}\]

OpenStudy (zarkon):

last inequality should be \(<\)

OpenStudy (anonymous):

oh wow thank you ! @Zarkon

OpenStudy (zarkon):

np